OpenStudy (applied):

What is the equation of the line which includes points (2, 3) and (0, 10)?

zarkam21 (zarkam21):

https://www.desmos.com/calculator Use this =)

OpenStudy (applied):

idk how to use that

OpenStudy (applied):

A. y equals negative seven halves x minus ten B. y equals negative two sevenths x plus ten C. y equals negative seven halves x plus ten D. y equals two sevenths x plus ten these are the options

OpenStudy (nuttyliaczar):

Find the slope between the two points

OpenStudy (nuttyliaczar):

That's the change in y over change in x

MARC (marc_d):

Step 1: Find the slope. Slope formula\[\huge{m=\frac{ y_2-y_1 }{ x_2-x_1 }}\]let say\[\huge{(2,3)=(x_1,y_1)}\]\[\huge{(0,10)=(x_2,y_2)}\]plug in all the values into the slope formula to find the slope,m.\[\huge{m=\frac{ 10-3 }{ 0-2 }}\]\[\huge{m=-\frac{ 7 }{ 2 }}\]Step 2: Use the slope-intercept form to find the value of c(y-intercept)\[\huge{y=mx+c}\]where m=slope c=y-intercept Then,choose any of those two points to represent as\[\huge{(x,y)}\]Let say, i choose point\[\huge{(2,3)=(x,y)}\]now,plug in all the values into the slope-intercept form(y=mx+c) to find the value of c(y-intercept)\[\huge{y=mx+c}\]\[\huge{3=2(-\frac{ 7 }{ 2 })+c}\]\[\huge{3=-7+c}\]\[\huge{c=10}\]Step 3: Form the equation of the straight line using y=mx+c\[\huge{y=mx+c}\]\[\huge{y=-\frac{ 7 }{ 2 }x+10}\] @Applied